OpenStudy (ahsome):

What is the equation of a circle, when the radius is 5 and the points pass through (3, 1) and (-2, -4)

OpenStudy (ahsome):

I know the equation is:\[(x-h)+(y-k)=r^2\] Substituting the radius, I get:\[(x-h)+(y-k)=25\] I am stuck there

OpenStudy (anonymous):

SO whas the p[roblem here....you 0nly need to plug-in the values?

OpenStudy (anonymous):

what coordinate you are taking as centre?

OpenStudy (ahsome):

The centre is where 5 points away from both (3,1) and (-2,-4)

OpenStudy (ahsome):

The answer is this, if that helps\[(x+2)^2+(y-1)^2=25\]

OpenStudy (anonymous):

very easy just take any one of the points as reference and u will get the centre point and then plug the values

OpenStudy (ahsome):

Sorry, but I am still not understanding using one of the points as reference? What do you mean?

OpenStudy (anonymous):

say you take the P(3,1) As reference and subtract 5 points from it such that twice the point will yield P(-2,-4)...got it?

OpenStudy (anonymous):

and then you can get the centre and plug the values then and you will get the answer nah...you have to make it such that twice the 5 gives you negative coordinates points..and here's the catch.. use the euclidean distance formula........ and construct it

OpenStudy (ahsome):

OK, I give up on this question. I don't even know what the euclidean distance formula is. But thanks anyway

OpenStudy (anonymous):

hehey..why give up... euclidean distance means the distance between 2 points in a Cartesian plane: given by \[\xi = \sqrt{(x_1 - x_2)^2 + (y_1 - y_2)^2}\] They should have taught it to u but haven't give you the 0fficial name

OpenStudy (anonymous):

@Ahsome :NEVER GiVE UP

OpenStudy (ahsome):

I am in Grade 10. So I don't know

OpenStudy (ahsome):

Ok, doing that will give me what number? The Radius?

OpenStudy (anonymous):

OK I have made it quite complex instead of using other point as reference(the negative coordinate ones(-2,-4)) just use the positive ones as reference ..and get the coordinates for the centre of the circle (5 [points away)and then use those values to plug in the formula and Voila..you will get the answer

OpenStudy (anonymous):

()K I gotta go..I will give you the workin' when I will com back if you will find yarself' unable to do it...just think.....otherwise the wonderful community of OS(including me ..lol) will be there...

OpenStudy (ahsome):

I am sorry for being so unreasonable, but I still am not sure. Can you like draw it or something?

OpenStudy (ahsome):

K, you go. Whenever you can, please help ;) @Adjax

OpenStudy (ahsome):

If you can help @shereenkhan ,please do

OpenStudy (anonymous):

i am trying bro :)

OpenStudy (ahsome):

Thanks man :D

OpenStudy (ahsome):

Come on @iambatman , instead of saving Gotham try to save my Maths B test ;)

OpenStudy (anonymous):

@aryandecoolest. we need help

OpenStudy (anonymous):

what formula u were using is for (h,k)---->is co ordinates of the center and radius,r=5. so it won't be applicable here :P

OpenStudy (anonymous):

OH man..... here's hint 5 = sqrt((x1 -x2) + (y1-y2)) put the values of x1 and y1 you will get the eqn. x +/- y = something(+/- denotes I'm not sure of what symbol will come after operation) repeat the step once more with the other point (-2,-4) and you willl get again x /- y = somethijng I assume you know about simultaneous EQNs. and solving them(Since you are 10th grader) => x +/- y = something x +/- y = something ------------------- ______________________ you will get the value of any one and using anyone you can find the other variable and voila..you'll get the answer

OpenStudy (anonymous):

remember the euclidean distance formula it's the key

OpenStudy (anonymous):

@Ahsome

OpenStudy (anonymous):

hello @Adjax its not the way to solve it.. it lil tricky

OpenStudy (anonymous):

two points on circle are given .. putting in (x-h)2+(y-k)2=r2 you have to putt these points and radius and get two equations in variables h and k so to solve them u have to find out the center of circle

OpenStudy (anonymous):

after finding center (h,k) u can find out equation of circle ..

OpenStudy (anonymous):

so radius value lena hai.

OpenStudy (anonymous):



OpenStudy (anonymous):

okay :) ous ke baath.

OpenStudy (anonymous):

look u hv to go in thsi direction (3-h)^2+(1-k)^2=5^2 .... eq1 (-2-h)^2+(-4-k)^2=5^2 .... eq2

OpenStudy (anonymous):

in do equations ko simultaneouslu solve kro h & k find kro .. mtlb center (h,k) then u can find equation by putting radius and center in (x-h)^2+(y-k)^2=r^2 where (h, k) is center and r is radius.. so tahts all Lol!

OpenStudy (anonymous):

that's all :)

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

sahe kahu mere upper se chala gaya :)

OpenStudy (anonymous):

i mean ser k upper se :)

OpenStudy (anonymous):

easy c bt hy .. 2 points pata hen circle pr .. or radius b .. ab is sb ki help se hm ne center find out krna hy .. itan to smjh lg ra hy na ??

OpenStudy (anonymous):

q k circle ki equation k lye hmen just 2 chezen pta honi chahye .. aik center (h,k) or aik radius ... and radius to provided hy lakin center ni hy .... to sb se phle socho k center kese find out kren

OpenStudy (anonymous):

h,k ke value lena hai. right:)

OpenStudy (anonymous):

u knw.. (x_h)^2_(y-k)^2=r^2 is me radius or 1st point put kren to u can have equation in form of h and k ... ryt issi terha radius or 2nd point putt krne se h or k me 2nd equation mil jye gi ..... ab in dono h k ki equations ko solve krne se hmen h or k ki values mil jyen gi

OpenStudy (anonymous):

yup ryt

OpenStudy (anonymous):

iss k baath i am blank.

OpenStudy (anonymous):

ab center find kr lo gi na h and k ???

OpenStudy (anonymous):

jb yeh find kr liya na to center ki equation (x-h)^2+(y-k)^2=r^2 h k ki wavlues or radius ki value put kr k equtaion nikal lena simple

OpenStudy (anonymous):

hw to find a center point

OpenStudy (anonymous):

arey yra .. h, k center he hy

OpenStudy (anonymous):

R(h,k) its center

OpenStudy (anonymous):

toh ous ka value kise nikalu.

OpenStudy (anonymous):

to oper mene phr se kia btaya hy woh smjho

OpenStudy (anonymous):

fist equ... ku 2equ... ku kiya karu.

OpenStudy (anonymous):

ku k equation 2 veriables me ho gi h and k ....... so h and k ikathe aik he equation me kese solve kro gi us k lye 2 equations zaroori hen ..

OpenStudy (anonymous):

koin sa yeh (-2-h)^2+(-4-k)^2=5^2

OpenStudy (anonymous):

@khurramshahzad : I can't get you correctly ..what language you are speaking(I used google but no language is detected)...BUt what I've got from yar working is quite right.. But my method is also not wrong .. As you already described your method is little tricky(not much but yes it is for the Asker) ...I just simplified my approach 'cuz I think that the @Ahsome is having a little trouble....

OpenStudy (anonymous):

yup is ko solve kro to h^2 and k^2 ki form me equation aye gi wahan se h ko seperate jrna

OpenStudy (anonymous):

and can you summarize all yar' working in one post..it will be better....

OpenStudy (anonymous):

mmm @Adjax did u see my 2nd comment .. it will tell u the direction of questions. i went through very simplified way

OpenStudy (anonymous):

I meant that a summarized post will l00k much better and will be easy to digest and I can't understand some of yar' post..I can only got some bits right...

OpenStudy (anonymous):

wht did u said @khurramshahzad

OpenStudy (anonymous):

and the method I described in some indirect forms thought very simple is used for the same purpose such as finding a distance between object in computer gaming....

OpenStudy (anonymous):

(-6-h)^2+(-4-k)^2=5^2

OpenStudy (anonymous):

two points on circle are given .. putting in (x-h)2+(y-k)2=r2 you have to putt these points and radius and get two equations in variables h and k so to solve them u have to find out the center of circle after finding center (h,k) u can find out equation of circle .. look u hv to go in thsi direction (3-h)^2+(1-k)^2=5^2 .... eq1 (-2-h)^2+(-4-k)^2=5^2 .... eq2 solve these 2 equations simultanously and find h and k taht will be center (h,k) then u can find equation by putting radius and center in (x-h)^2+(y-k)^2=r^2 where (h, k) is center and r is radius.. so tahts all Lol! u hv two points on circle .. and radius .. now find out center h and k thats center .. then using center and radius you will hv teh required equation

OpenStudy (anonymous):

okay @Adjax may be u are right too.. actually i did not get ur meanings

OpenStudy (anonymous):

anyway...there are 100 ways to get to destiny....l0lz ./.:+UP

OpenStudy (anonymous):

lol!

OpenStudy (anonymous):

and that language is urdu .. i thought we all were urdu speaking on this qeuestion :D